OpenStudy (anonymous):

Find an equation in slope-intercept form of the line satisfying the specified conditions? Through ( 8 , - 5 ), parallel to 8 x + 3 y = 43

OpenStudy (kc_kennylau):

Find out the slope of the line first? :)

OpenStudy (anonymous):

I dont know how to dot his. I put it on yahoo answers and the person who answer got the same answer as I did..and we are both wrong acccording to my homework lol

OpenStudy (kc_kennylau):

what have you got?

OpenStudy (anonymous):

Sorry for not replying quickly my computer is being inredibly annoying tthe past few days. It's like it knows Im studying for finals ;_;

OpenStudy (kc_kennylau):

nah OS was down

OpenStudy (kc_kennylau):

maybe caused by server overloading coz as u said it's the finals

OpenStudy (anonymous):

(y+5) = (-8/3)*(x-8) or y = -(8/3)x +(24/3) - 5 = -(8/3)x +9 or] finally y = -(8/3)*x + 9

OpenStudy (anonymous):

Oh! You are probably right along with my blackboard being sluggish, too.

OpenStudy (anonymous):

So even though y in the (8,-5) is negative when I put it in for y+5= -8/3(x-8) I turn it into a positive??

OpenStudy (kc_kennylau):

?

OpenStudy (anonymous):

On the wolf website it shows it as positive even though the 5 in the parenthesis is negative